A pendulum that has a period of 7.01 s and that is located where the acceleration due to gravity is 9.44 m/s2 is moved to a location where the acceleration due to gravity is 9.31 m/s2. What is its new period?



Round your answer to 5 decimal places.



Pls dont forget the unit

1
Expert's answer
2022-05-14T16:45:44-0400
"T_1=2\pi\sqrt{\frac{l}{g_1}}"

"T_2=2\pi\sqrt{\frac{l}{g_2}}"

"T_2=T_1\sqrt{\frac{g_1}{g_2}}=7.01\sqrt{\frac{9.44}{9.31}}=7.06\:\rm s"